OpenStudy (anonymous):

Solve the following equation for p: q = 4p - 11

OpenStudy (ajprincess):

in order to solve for p u have to isolate p. see p is not alone. it has been multiplied by 4 and 11 has been subtracted from it. so adding 11 to 4p-11 will give u 4p. bcos +11-11=0 4p+0=4p if u add 11 to right side of the equation then u have to add 11 to left side too to make the equation valid. so the left side of the equation will be q+11. nw u will have q+11=4p getting this? @BeatriceM2012

OpenStudy (anonymous):

Ok; first we could have 4p=q+11 so now divide the sides by 4 you will have p=1/4 q +11/4

OpenStudy (anonymous):

@ajprincess - Yeah. I get it.

OpenStudy (ajprincess):

good:) nw u see that u have 4p on the right side. but u need to have only p. so vat do u need to do?

OpenStudy (anonymous):

Divide each side by 4.

OpenStudy (ajprincess):

ya good:) so vat will be the final answer?

OpenStudy (anonymous):

1/4q + 11/4 = p

OpenStudy (ajprincess):

yup:)